how long does it take a naturally growing bacteria population to quadruple if it doubles in 6.82 hours?
Answer by ankor@dixie-net.com(22740)

how long does it take a naturally growing bacteria population to quadruple
if it doubles in 6.82 hours?
:
The doubling time growth model
Ao*2^(t/d) = A
Where
A = resulting amt (4)
Ao = initial amt (1)
d = doubling time (6.82)
t = time
:
1*2^(t/6.82) = 4
We can write it:
2^(t/6.82) = 2^2
Base is equal, therefore the exponents are
t/6.82 = 2
t = 2 * 6.82
t = 13.64 hr to quadruple
: